# Quantum Computing Education in Latin America: Experiences and Strategies

### v1

@@ II. QT Hardware and Software Engineering (Target level: A1 - Awareness)
  • Example: “What is a qubit, and how does it differ from a classical bit? (a) A qubit can represent both 0 and 1 simultaneously (b) A qubit follows classical computing rules.”
@@ II-C Future vision
Building on the foundation of introductory and basic QC courses, we aim to develop further undergraduate and graduate specialization courses to gain deeper understanding and practical skills in specific areas. The courses would focus on achieving B and C proficiency levels and qualification profiles from P5 to P9 depending on students’ profile, interests and experience. Furthermore, as QT continue to evolve rapidly, it is imperative to establish continuous professional development programs for educators, providing ongoing training and resources to keep instructors abreast of the latest advancements. In addition, training courses based on the ECFQT can be offered to students with more advanced knowledge, enabling them to become educators themselves. This approach not only helps increase the number of trained faculty but also facilitates the dissemination of knowledge to a broader student base, thereby expanding the reach and impact of QC education.
@@ III Results and Discussion
In addressing the challenge of quantum education and training, we aimed to close the gap in Latin America and the rest of the world with this first step of course creation, contributing to the unification of scientific and technological progress. We highlighted the contributions of the UNAL’s research group, an initiative that can serve as a model for other groups aiming to contribute to the field and its expansion. We also provided a practical and illustrative example of how to implement and adapt the European Competence Framework for Quantum Technologies, which not only guided the profiles and proficiency levels needed but also offered valuable information on needs, suggestions, and learning paths for each profile, useful to students once they end the course. Additionally, we introduced the new elements of “Work Teams” and “Evaluation” based on the framework to enhance the courses’ development. These additions were made possible thanks to the framework’s clarity and comprehensibility, as it offers numerous tools that allow educators to maximize its potential and create additional, valuable elements for education. Further advancement of these tools will involve examining and creating additional examples and methodologies. Once completed, these could potentially be integrated as official sections within the framework, as their significance lies in illustrating the learning’s applicability and unifying assessment methods, thereby contributing to a more robust and standardized approach to QC education.
@@ IV Conclusions
One of the key challenges addressed is accessibility, both in terms of geographic reach and inclusivity for students lacking enough foundations in the field. By enhancing access to QC education across Latin America, we can leverage the diverse talent pool in Latin America, fostering inclusivity and a new generation of professionals who bring varied perspectives and innovative solutions to the global stage. This will not only elevate the region’s contribution to the global quantum ecosystem but also help bridge the growing skills gap worldwide, driving forward advancements in technology and its applications across various sectors. While the courses represent a step forward, further efforts are needed to ensure their scalability and adaptability. Potential limitations include relevant content selection and logistical challenges of delivering high-quality, practical QC education through online platforms.

### v2

@@ II. QT Hardware and Software Engineering (Target level: A1 - Awareness)
  • Example: “What is a qubit, and how does it differ from a classical bit? (a) A qubit can exist in a superposition of both 0 and 1 states. (b) A qubit follows classical computing rules.”
@@ IV. Overall Evaluation Proposal:
The aim is to ensure a balanced evaluation of both theoretical understanding and practical application, in alignment with the framework’s competencies. Slightly more weight is given to the first area, as the goal is to ensure students have a strong theoretical foundation in quantum concepts. This will provide them with the necessary background to better focus on software, hardware, and applications in subsequent courses. Therefore, in each of the three areas, a greater emphasis has been placed on knowledge assessment rather than skills assessment.
@@ II-C Future vision
Building on the foundation of introductory and basic QC courses, we aim to develop further undergraduate and graduate specialization courses to gain deeper understanding and practical skills in specific areas. The courses would focus on achieving the B proficiency level and qualification profiles from P4 to P6 depending on students’ profile, interests and experience; as reaching the C level and the P7, P8 and P9 profiles typically requires extensive work experience or advanced research, such as a PhD or post-doc. Furthermore, as QT continue to evolve rapidly, it is imperative to establish continuous professional development programs for educators, providing ongoing training and resources to keep instructors abreast of the latest advancements. In addition, training courses based on the CFQT can be offered to students with more advanced knowledge, enabling them to become educators themselves. This approach not only helps increase the number of trained faculty but also facilitates the dissemination of knowledge to a broader student base, thereby expanding the reach and impact of QC education.
@@ III Results and Discussion
In addressing the challenge of quantum education and training, we aimed to close the gap in Latin America and the rest of the world with this first step of course creation, contributing to the unification of scientific and technological progress. We highlighted the contributions of the UNAL’s research group, an initiative that can serve as a model for other groups aiming to contribute to the field and its expansion. We also provided a practical and illustrative example of how to implement and adapt the European Competence Framework for Quantum Technologies, which not only guided the profiles and proficiency levels needed but also offered valuable information on needs, suggestions, and learning paths for each profile, useful to students once they end the course. Additionally, we introduced the new elements of “Work Teams” and “Evaluation” based on the framework to enhance the courses’ development. These additions were made possible thanks to the framework’s clarity and comprehensibility, as it offers numerous tools that allow educators to maximize its potential and create additional, valuable elements for education. Further advancement of these tools will involve examining and creating additional examples and methodologies. Once completed, these could potentially be included in a work showcasing the applications of the CFQT, as their significance lies in illustrating the learning’s applicability and unifying assessment methods, thereby contributing to a more robust and standardized approach to QC education.
@@ IV Conclusions
One of the key challenges addressed is accessibility, both in terms of geographic reach and inclusivity for students lacking enough foundations in the field. By enhancing access to QC education across Latin America, we can leverage its diverse talent pool, fostering inclusivity and a new generation of professionals who bring varied perspectives and innovative solutions to the global stage. This will not only elevate the region’s contribution to the global quantum ecosystem but also help bridge the growing skills gap worldwide, driving forward advancements in technology and its applications across various sectors. While the courses represent a step forward, further efforts are needed to ensure their scalability and adaptability. Potential limitations include relevant content selection and logistical challenges of delivering high-quality, practical QC education through online platforms.